3x
--- - 1/3 = 2
2

Use cross-multiplication to solve.
How do I get rid of the 2 to use cross-multiplication? Non-homework.

Found 2 solutions by rothauserc, Boreal:
Answer by rothauserc(4718)

(3x/2) - (1/3) = 2
:
we use 6 which is a common multiple of 3 and 2
:
(3(3x)/6) -2/6 = 12/6
:
(9x/6) -(2/6) = 12/6
:
9x/6 = 14/6
:
cross multiply fractions
:
54x = 84
:
x = 84/54 = 14/9
:
check answer
:
(3(14/9)/2) - (1/3) = 2
:
(42/18) - (6/18) = 2
:
36/18 = 2
:
2 = 2
:
answer checks
:

Answer by Boreal(15235)

3x/2-(1/3)=2
3x/2=2+(1/3)=(7/3); I added 1/3 to both sides. 2= 6/3, so 2 + (1/3)=7/3.
now cross-multiply
9x=14
x=14/9
check
3(14/9)/2-1/3=42/9 divided by 2 -1/3
That is 42/18-1/3
that is 42/18-6/18=36/18=2